# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Abstract
A constitutive relation between stress and strain relative to a reference state is the basic assumption of elasticity theory. However, in living matter, stress is governed by (motor molecule) activity rather than a constitutive law. What paradigm takes the place of elasticity in this setting? Here, we derive a continuum theory of active mechanics by taking the continuum limit of the Active Tension Network model of 2d epithelia. Instead of a reference state, we start from a prescribed active force configuration, encoded in a Riemannian “tension metric”. Intuitively, one expects cells to adjust their positions to achieve force balance by rearranging local sources of active stress. More precisely, the cell positions define an embedding of the tension metric into 2d physical space, which determines the macroscopic physical stress. For free boundaries, tissue adopts a certain intrinsically defined shape, the force-balanced embedding with minimal internal stress. Boundary forces then deform this embedding. The resulting stress transformation yields an effective stress-strain relation. Key elements of elasticity hence emerge from a “stress-only” starting point, explaining how tissue shape can be adiabatically controlled by active stress during morphogenesis. Plastic behavior arises from topological cell rearrangement, which we represent by a continuous reparameterization of the tension metric, providing a principled continuum theory of emergent elasto-plastic flow. To express this physics, we use the mathematics of isothermal coordinates and quasi-conformal maps. An explicit coarse-graining study, described in a companion paper, substantiates the results of our continuum analysis. The present theory elucidates the unconventional mechanics of living tissues and may apply to 2d active and granular materials more generally.
@@ Introduction
Conventional elasticity, based on a reference shape and a constitutive law relating stress and strain (rate) [3], has served as the starting point for much theoretical work on tissue mechanics [4, 5, 6]. For instance, epithelial cells are often described as having a target area and perimeter [7] with deformation from a target shape defining the stress. Tissue flow on longer time scales is understood in terms of plastic change of the reference state attributed to cell division or rearrangement  [2, 8].
@@
However, biological observations challenge the notions of reference state and constitutive laws. First, the turnover of the force-bearing molecules in cells, rapid compared to morphogenetic timescales, allows the “rest length” of a cell to change [9], ruling out a fixed reference state. Since stress needs to be actively maintained against relaxation due to turnover, one expects that on long timescales, all stress in a tissue is active. Active stress, however, is in principle independent of strain and instead controlled by the activity of molecular motors such as myosin.
@@
We therefore seek to understand tissue mechanics by starting from the specification of cell-associated active stress, instead of a reference shape. Due to the timescale separation between mechanical equilibration (seconds [10]) and morphogenetic remodeling (hours), we expect cells to adjust their shapes and positions to achieve local force balance. Morphogenesis takes the form of an adiabatic transformation between force-balanced states, controlled by changes in the local mechanical activity.
@@
This approach is motivated by the study of confluent, 2D epithelia. Cells in such tissue sheets can regulate their interfacial tensions by molecular motor activity in a cell’s actomyosin cytoskeleton [11, 12], which acts as a “microscopic muscle”. This picture is made precise by the active tension network (ATN) model [11], a minimal model for the mechanics of 2D epithelia. Within this framework, we can ask how cells control tissue shape through local active tensions.
@@
To build intuition, consider a conventional fluid foam. Here, a “cell” (foam bubble) interface has no rest length since its tension is independent of its length, being instead set by the fluid’s surface tension. These surface tensions are balanced against internal pressure in the foam bubbles. Remarkably, even though microscopically fluid, the foam macroscopically behaves like an elastic solid – it can maintain its shape against gravity [13]. In a foam, interfacial tensions are fixed material parameters. In a tissue, they are promoted to active, dynamical variables (under local cellular control), enabling richer behavior.
@@
Here, we derive a continuum theory of emergent, active elasticity. The key to our approach lies in representing the configuration of active stress by a Riemannian metric $\mathbf{g}$, a continuum version of Maxwell–Cremona force tesselations [14]. The cell tiling defines an embedding of the tension metric into physical space. We show that the central elements of elasticity theory, a stress-free reference state and a stress-strain relationship, emerge from this embedding. Additionally, we will demonstrate that cell-scale topological rearrangements that enable large deformations can be captured in the continuum theory via a suitable reparametrization of the tension metric. Taken together, these ideas enable a continuum description of adiabatic tissue flow driven by active stress dynamics. The resulting theory is a form of “emergent” viscoelasticity, in which adiabatic stress dynamics can drive active plastic flow in the absence, or against the resistance, of external forces. In a companion paper [15], we show that the continuum-based, “top-down” approach presented here precisely matches a “bottom-up” coarse-graining analysis of the ATN model.
@@
The remainder of this paper is structured as follows. We begin by reviewing the cell-level ATN model. In Section .1, we explain how we describe cell tiling and tension configuration in the continuum limit. Section .2 studies the static problem, and Section .3 turns to the adiabatic dynamics of active stress. Finally, Sect. .4 considers topological cell rearrangement.
Discrete theory: 2D tension networks in force balance
@@
The microscopic ATN model [11] describes a 2D tiling of cells $i,j,\dots$ whose mechanics is dominated by tensions $\tau_{ij}$ along cell-cell interfaces and intracellular pressure $p_{i}$. In mechanical balance, virtual work has to vanish, $0=dE=-\sum_{ij}\uptau_{ij}d\ell_{ij}+\sum_{i}\mathrm{p}_{i}d\mathrm{A}_{i}$, where $\ell_{ij},\mathrm{A}_{i}$ are interface lengths and cell areas. The pressures $\mathrm{p}_{i}$ are determined by passive bulk elasticity, captured by a local equation of state $\mathrm{p}_{i}=P(\mathrm{A}_{i})$. By contrast, the tensions $\tau_{ij}$ are dynamical variables determined by intrinsic motor molecule activity, and independent of interface length $\ell_{ij}$ – there are no reference lengths and no constitutive laws for tensions. This is the key premise of the ATN model [11, 16].
@@
We aim to understand how the active tensions determine the physical configuration of the cell tiling in force balance, and how this configuration evolves when tensions change adiabatically. To this end, it will be useful to think of the tension configuration as a network dual to the cell tiling. Since the cells generically meet at 3-fold vertices, the tensions form a triangulation with one node per cell $i$ and a length $\uptau_{ij}$ link between neighbors $i$ and $j$; Fig. 1. Each tricellular vertex ijkijk corresponds to one “tension triangle”, the shape of which encodes the local tension configuration (for example, tension anisotropy).
@@ Results
.1 Continuum description of ATNs as Riemannian surfaces
@@
A triangulation, specified by its adjacency graph and edge lengths, defines a (piecewise-linear) metric surface  [^1] The definition of the manifold structure is explained in detail in our companion paper [15]. (see Fig. 1, top). We can therefore coarse-grain the tension triangulation to a continuous tension manifold with tension metric $g_{\alpha\beta}(\bm{\xi})$ as a function of Lagrangian coordinates $\bm{\xi}$, which play the role of the cell labels $i,j$ For an infinitesimal line element $d\bm{\xi}$, the tension metric yields the tension $d\tau$ on cell interfaces oriented transverse to it as $d\tau^{2}=g_{\alpha\beta}(\bm{\xi})d\xi_{\alpha}d\xi_{\beta}$. In the “microscopic limit” where $d\bm{\xi}$ connects two adjacent triangulation nodes, $d\tau$ is the tension along the interface between the corresponding cells.
Note that different triangulations (i.e., local tension configurations) can realize the same tension manifold. For instance, all triangulations of the plane realize the Euclidean metric. The tension triangulation thus encodes “microscopic” information in addition to the tension metric $\mathbf{g}$; we return to this point systematically in Sect. .4. One important additional degree of freedom is the local cell density $n({\bm{\xi}})d^{2}\xi$, the number of cells (triangulation nodes) per unit area.
A set of physical cell positions $\mathbf{r}({\bm{\xi}})$ now defines a map from the tension manifold into two-dimensional physical space, to which we refer as “embedding” (Fig. 1, bottom). For simplicity, we assume that physical space is the plane, but our theory readily generalises to tissues on arbitrary curved surfaces. Under the map $\mathbf{r}({\bm{\xi}})$, the metric $\mathbf{g}$ transforms as
@@
We now link the geometry of the tension metric to the mechanics of the cell. The macroscopic stress tensor of an ATN is the sum of two contributions
@@
where $\sigma_{\alpha\beta}$ is the tensional stress due to tensions $\uptau_{ij}$, and $p$ the intracellular pressure. The original ATN model [11] assumed uniform intracellular pressure. Here, we generalize: intracellular pressure is determined by an equation of state $p=P(n)$ as a function of cell density in the continuum. The reference cell density $n_{0}$ is defined by $P(n_{0})=p_{0}$. (In contrast to the active tensile stress $\sigma$, we treat pressure as a passive field, hence the equation of state.) Under $\mathbf{r}({\bm{\xi}})$, the density transforms as $n(\mathbf{r})=n({\bm{\xi}})\,{\det[\partial_{\bm{\xi}}\mathbf{r}]}$, and therefore:
In the following, we refer to $p$ simply as pressure, noting that it differs from the total pressure $\mathrm{Tr}[\bm{\sigma}^{\mathrm{tot}}]/2=p-\mathrm{Tr}[\bm{\sigma}]/2$.
The tensile stress $\bm{\sigma}$ is determined by the tension metric. Consider two adjacent cells at $\mathbf{r}$ and $\mathbf{r}+d\mathbf{r}$ (we therefore implicitly set $dr$ to be of the scale of a single cell, which becomes infinitesimally small in the continuum limit). The magnitude of the force $d\bm{f}$ acting through the virtual cut $d\mathbf{r}$ (Fig. 2) equals the tension $|d\bm{f}|=d\tau$ along the interface between the two cells. It is determined by the tension metric, $d\tau^{2}=p_{0}^{2}\,g_{\alpha\beta}(\mathbf{r})dr_{\alpha}dr_{\beta}$. On the other hand, the stress tensor applied to the cut normal yields $d\bm{f}=\sigma\cdot\mathbf{n}\cdot dr=\mathbf{\sigma}\cdot\epsilon\cdot d\mathbf{r}$, where $\epsilon$ is the antisymmetric tensor, $\mathbf{n}$ is the unit normal, and $\cdot$ is the standard inner product in Euclidean physical space. Equating the tension- and cell-space expressions:
$p_{0}^{2}d\mathbf{r}^{T}\cdot\mathbf{g}(\mathbf{r})\cdot d\mathbf{r}=d\tau^{2}=d\mathbf{r}^{T}\cdot(\epsilon^{T}\cdot\bm{\sigma}^{2}(\mathbf{r})\cdot\epsilon)\cdot d\mathbf{r}$ (4)
Since this equation must hold for arbitrary $d\mathbf{r}$, the stress tensor must be the matrix square root of the metric, rotated by $\pi/2$, and scaled by $p_{0}$. This relation between microscopic stress, encoded in the tension metric, and macroscopic stress is a central element of our continuum theory. Eq. (4) and the stress tensor are defined in physical space, not on the tension manifold, and depend on the cell positions $\mathbf{r}({\bm{\xi}})$, which are physical observables, not mere coordinates.
In mechanical equilibrium, forces must balance:
@@
Equations (4) and (5), supplemented with mechanical boundary conditions (BCs), set the physical configuration (cell positions) $\mathbf{r}({\bm{\xi}})$ for a given tension metric $\mathbf{g}({\bm{\xi}})$ and cell density $n({\bm{\xi}})$. In the following, we introduce a mathematical framework to solve Eqs. (4)–(5). We first study the statics problem to see how tissue shape and response to external forces emerge from (fixed) $\mathbf{g}({\bm{\xi}})$. Then, we turn to quasi-static shape dynamics due to changes of $\mathbf{g}({\bm{\xi}})$, and finally generalize our framework to describe plastic flow (cell rearrangement).
Figure 2: Left: In the continuum limit, the displacement vector between two cell centroids becomes a differential $d\mathbf{r}$. The tension metric determines the tension $d\tau$ on the interface between these cells. In isothermal coordinates, $dr=p_{0}^{-1}d\tau$. Right: The stress tensor $\mathbf{\sigma}$ is defined by the traction forces that act through infinitesimal virtual cuts. A cut along $d\mathbf{r}$ must yield the interfacial tension, relating tension metric and stress tensor in Eq. (4).
@@
To solve Eq. (5), we use complex coordinates $z=z_{1}+iz_{2}$. A symmetric matrix $\mathbf{q}$ (e.g. the metric $\mathbf{g}$) defines a quadratic form $\mathbf{v}\mapsto q_{\alpha\beta}v^{\alpha}v^{\beta}$ which becomes
$4q_{\alpha\beta}v^{\alpha}v^{\beta}$ $=(q_{z\bar{z}}+q_{\bar{z}z})v\bar{v}+q_{zz}vv+q_{\bar{z}\bar{z}}\bar{v}\bar{v},$ (6a)
@@
Since $q_{zz}$ is the deviatoric component of $\mathbf{q}$, the Beltrami coefficient $\mu_{q}$ measures the relative anisotropy
$\mu_{q}=q_{zz}/\mathrm{Tr}[\sqrt{\mathbf{q}}]^{2}$ (7)
@@
Maps satisfying the Cauchy–Riemann equation $\bar{\partial}_{z}f=0$, i.e. $z\mapsto z^{\prime}=f(z)$, are conformal (angle-preserving). Their Jacobian is a diagonal matrix ${\mathbf{D}\kern-0.80002ptf}=\mathrm{diag}[\partial_{z}f,\overline{\partial_{z}f}]$, combining rotation and scaling by the conformal factor $\lambda_{z|z^{\prime}}=|\partial_{z}f|$. Furthermore, in this case, $\Delta\log\lambda_{z|z^{\prime}}=4\partial_{z}\bar{\partial}_{z}\operatorname{Re}[\log(\partial_{z}f)]=0$. More generally, the shear created by a map is measured by its Beltrami coefficient $\mu_{z|z^{\prime}}=\bar{\partial}_{z}f/\partial_{z}f$ (the subscript of $\mu_{z|z^{\prime}}$ denotes the map’s source and target.) Note that the Beltrami coefficients of the map $f$ and the quadratic form $({\mathbf{D}\kern-0.80002ptf})^{\dagger}\cdot{\mathbf{D}\kern-0.80002ptf}$ agree; $()^{\dagger}$ being the Hermitian adjoint. Geometrically, $f$ maps infinitesimal circles to ellipses with eccentricity $|\mu_{z|z^{\prime}}|/(1-|\mu_{z|z^{\prime}}|^{2})^{1/2}$. If $|\mu_{z|z^{\prime}}|<1$ everywhere, $f$ is called quasi-conformal (QC). In particular, QC maps preserve orientation. Notation-wise, their argument distinguishes QC maps $f(z,\bar{z})$ (also written $f(\mathbf{z})$), from conformal maps $f(z)$. In the following, we use the formalism of QC maps to develop our continuum theory (in Ref. [18], a similar QC approach was used to describe the growth of 2d tissues).
@@
$4g_{\alpha\beta}d\xi^{a}d\xi^{b}$ $=\mathrm{Tr}[\sqrt{\mathbf{g}}]^{2}|d\xi+\mu_{g}d\bar{\xi}|^{2}$ (9)
@@
where the metric’s Beltrami coefficient $\mu_{g}=g_{zz}/\mathrm{Tr}[\sqrt{\mathbf{g}}]^{2}$ measures the anisotropy of $\mathbf{g}$. Therefore, $z(\xi,\bar{\xi})$ must fulfill a Beltrami equation:
$\bar{\partial}_{\xi}z(\xi,\bar{\xi})=\mu_{g}\partial_{\xi}z(\xi,\bar{\xi})$ (10)
(Equivalently, $\mu_{\xi|z}=\mu_{g}$.) The QC map from Lagrangian $\xi$- to isothermal $z$-coordinates is exactly such that it removes the shear part of the metric. (By abuse of notation, we will denote both the coordinates and the map $\xi\mapsto z(\xi,\bar{\xi})$ by $z$). The metric, therefore, becomes isotropic and is fully described by its conformal factor $\lambda_{g}$
@@
Isothermal coordinates are conformal and do not distort angles (see Fig. 3). Appendix G.2 discusses an example (isothermal coordinates for a hemisphere).
In isothermal coordinates, the Liouville equation determines the Gaussian curvature:

### v3

# Emergent Elasticity and Quasiconformal Flow in Active Solids

@@ Abstract
A constitutive relation between stress and strain relative to a reference state is the basic assumption of elasticity theory. However, in living matter, force generation is governed by motor molecule activity, which does not depend on deformation relative to a reference. A different approach is needed to describe how cells sculpt tissues through local active forces. We develop a theory of two-dimensional continuum mechanics where the active stress configuration, rather than a reference shape, is the fundamental input. Motivated by the Active Tension Network model for epithelia, we encode motor-driven forces between cells in a Riemannian tension metric. We derive a stress–metric relation for the macroscopic stress that results from embedding the tension manifold into physical space (defining cell positions). Despite the absence of constitutive laws, a stress-free reference state and an effective stress-strain relation arise from the tension metric, making the system an effectively elastic active solid. Moving from statics to dynamics, our framework describes how an active solid can morph its shape through adiabatic dynamics of active stress. To capture large, plastic deformations through cell rearrangement, we introduce a second metric that geometrizes the cell network topology. Topological rearrangement appears as a continuous reparametrization of the tension manifold. This mathematical framework, based on Riemannian geometry, isothermal coordinates, and quasi-conformal flows, quantitatively predicts how local contractile activity determines large-scale shape and provides a principled continuum description of active plasticity. A companion paper validates the continuum analysis through coarse-graining of discrete cell networks. Our theory identifies a geometric origin of emergent elasticity and plasticity in living matter and, more broadly, in active and granular materials.
@@ Introduction
Conventional elasticity, based on a reference shape and a constitutive law relating stress and strain (rate) [3], has served as the starting point for much work on tissue mechanics [4, 5, 6]. For instance, epithelial cells are often described as having a target area and perimeter [7] with deformation from a target shape defining the stress. However, biological observations challenge this paradigm. First, the turnover of the force-bearing molecules in cells, rapid compared to morphogenetic timescales, allows the “rest length” of a cell to change [8], ruling out a fixed reference state. Since stress needs to be actively maintained against relaxation due to turnover, one expects active stresses to dominate on long timescales. Active stress, however, is not controlled by strain but by the activity of molecular motors, such as myosin. Nevertheless, tissues can maintain a macroscopic shape against external shear forces, and, more remarkably, change shape by internally controlled cell rearrangement during morphogenesis. We refer to such a system, which has neither a reference state nor a constitutive law, as an active solid.
@@
We therefore seek to understand tissue mechanics by starting from the specification of cell-associated active stress instead of a reference shape. The timescale of mechanical equilibration (seconds [9]) is well separated from that of morphogenetic remodeling (hours). One thus expects cells to adjust their shapes and positions to achieve force balance by rearranging local sources of active stress. Morphogenesis takes the form of an adiabatic transformation between force-balanced states, controlled by changes in the local mechanical activity.
@@
More precisely, we are motivated by the study of confluent, two-dimensional epithelia. Here, the dominant, tensile forces are generated along apical adherens junctions (cell-cell interfaces), which act as microscopic muscles [10, 11]. This picture is made precise by the active tension network (ATN) model [10], a minimal model for epithelial mechanics. Within this framework, we can ask how cells control tissue shape through local active tensions.
@@
To build intuition, consider a conventional fluid foam. Here, a “cell” (foam bubble) interface has no rest length since its tension is independent of its length, set instead by the fluid’s surface tension. These surface tensions are balanced against internal pressure in the foam bubbles. Remarkably, even though microscopically fluid, the foam macroscopically behaves like an elastic solid – it can maintain its shape against gravity [12]. In a foam, interfacial tensions are fixed material parameters. In a tissue, they are promoted to active, dynamical variables (under local cellular control), enabling richer behavior.
@@
We will represent the configuration of tensions by a Riemannian metric $\mathbf{g}$, a continuum version of Maxwell–Cremona force tesselations [13, 14]. The cell tiling defines an embedding of the tension metric into physical space. We show that the central elements of elasticity theory, a stress-free reference state and a stress-strain relationship, emerge from this embedding. Additionally, we will demonstrate that cell-scale topological rearrangements that enable large deformations can be captured in the continuum theory via a suitable reparametrization of the tension metric. Taken together, these ideas enable a continuum description of adiabatic tissue flow driven by active stress dynamics. The resulting theory is distinct from existing, active fluid models for morphogenesis [15, 16, 17], and makes differing, experimentally verifiable predictions for mechanical stress and the effect of motor molecule activity. In a companion paper [18], we show that the continuum-based, “top-down” approach presented here precisely matches a “bottom-up” coarse-graining analysis of the ATN model.
@@
The remainder of this paper is structured as follows. We begin by briefly reviewing the cell-level ATN model. In Section .1, we explain how we describe cell tiling and tension configuration in the continuum limit. Section .2 studies the static problem, and Section .3 turns to the adiabatic dynamics of active stress. Finally, Sect. .4 considers topological cell rearrangement.
Discrete theory: 2d tension networks
@@
The microscopic ATN model [10] describes a 2d tiling of cells $i,j,\dots$ whose mechanics is dominated by tensions $\tau_{ij}$ along cell-cell interfaces and intracellular pressure $p_{i}$ (Fig 1, bottom). Accordingly, the differential of elastic energy is given by $dE=-\sum_{ij}\uptau_{ij}d\ell_{ij}+\sum_{i}\mathrm{p}_{i}d\mathrm{A}_{i}$, where $\ell_{ij},\mathrm{A}_{i}$ are interface lengths and cell areas. In mechanical balance, $dE=0$. We assume that the cells’ positions $\mathbf{r}_{i}$ are constrained to a fixed two-dimensional surface, and that the cells do not exert traction forces on a substrate (therefore, the tissue-internal forces balance among themselves). This picture applies, for example, to the early Drosophila and chicken embryonic epithelia, which sit atop a liquid yolk and are constrained in the third dimension by the egg shell [1]. The pressures $\mathrm{p}_{i}$ are determined by a local equation of state $\mathrm{p}_{i}=P(\mathrm{A}_{i})$. (The original ATN model [10] assumed uniform $\mathrm{p}_{i}$). By contrast, the tensions $\uptau_{ij}$ are dynamical variables determined by intrinsic motor molecule activity. This is the key premise of the ATN model [10, 19].
Tissue dynamics unfolds on three distinct time scales. The speed of relaxation to mechanical equilibrium is the most rapid, on the order of seconds [9]. Turnover of force-bearing proteins within a few minutes sets the speed of relaxation of passive elastic stress in the cytoskeleton [20]. Finally, morphogenetic remodeling and regulation of contractile activity take place over tens of minutes to hours. Here, we exploit this hierarchy and consider only the dynamics at the slowest, morphogenetic timescale. We therefore assume that the ATN model is in instantaneous mechanical equilibrium and that transient passive stresses relax rapidly. In this limit, the interfacial tensions $\tau_{ij}$ are purely determined by local motor activity and independent of interface lengths $\ell_{ij}$ – there are no reference lengths and no constitutive laws for tensions.
@@
In this stress-first setting, we aim to understand how active tensions determine the physical configuration of the cell tiling in force balance, and how this configuration evolves when tensions change adiabatically. To this end, it will be useful to think of the tension configuration as a network dual to the cell tiling. Since the cells generically meet at 3-fold vertices, the tensions form an intrinsically defined triangulation with one node per cell $i$ and a length $\uptau_{ij}$ edge between neighbors $i$ and $j$; Fig. 1. Each tricellular vertex ijkijk corresponds to one “tension triangle”, the shape of which encodes the local tension configuration (for example, tension anisotropy). This Maxwell-Cremona-like “tension triangulation” serves as a point of departure for our analysis.
@@ Results
.1 Continuum description of ATNs by Riemannian geometry
@@
A triangulation, intrinsically specified by its adjacency graph and edge lengths, defines a (piecewise-linear) metric surface (see Fig. 1, top). We can therefore coarse-grain the tension triangulation to a continuous tension manifold with tension metric $g_{\alpha\beta}(\bm{\xi})$ as a function of Lagrangian coordinates $\bm{\xi}$, which play the role of the cell labels $i,j$. This continuum limit is well-defined as long as the microscopic tensions form non-degenerate triangles  [^1] The definition of the manifold structure and the continuum limit are explained in detail in our companion paper [18]. If tensions at a vertex violate the triangle inequality, this construction is not possible; however, as we will see below, such configurations are rapidly resolved by topological cell rearrangement.. For an infinitesimal line element $d\bm{\xi}$, the tension metric yields the tension $d\tau$ on cell interfaces oriented transverse to it as $d\tau^{2}=g_{\alpha\beta}(\bm{\xi})d\xi_{\alpha}d\xi_{\beta}$. In the “microscopic limit” where $d\bm{\xi}$ connects two adjacent triangulation nodes, $d\tau$ is the tension along the interface between the corresponding cells.
A set of physical cell positions $\mathbf{r}({\bm{\xi}})$ now defines a map from the tension manifold into two-dimensional physical space, to which we refer as the “embedding” (Fig. 1, bottom). For simplicity, we assume that physical space is the plane, but our theory readily generalises to tissues on arbitrary curved surfaces. Under the map $\mathbf{r}({\bm{\xi}})$, the tension metric $\mathbf{g}$ transforms as
@@
Note that different triangulations (i.e., local tension configurations) can realize the same tension manifold. For instance, all triangulations of the plane realize the Euclidean metric. The tension triangulation thus encodes “microscopic” information in addition to the tension metric $\mathbf{g}$; we return to this point systematically in Sect. .4. One important additional degree of freedom is the local cell density $n({\bm{\xi}})d^{2}\xi$, the number of cells (triangulation nodes) per unit area. The cell density determines the continuum pressure field $p$ (the equivalent of the intracellular pressures $p_{i}$) via an equation of state $p=P(n)$. In contrast to the active tensions, we hence treat pressure as a passive field. We define a reference cell density $n_{0}$ by $P(n_{0})=p_{0}$. Under $\mathbf{r}({\bm{\xi}})$, the density transforms as $n(\mathbf{r})=n({\bm{\xi}})\,{\det[\partial_{\bm{\xi}}\mathbf{r}]}$, and therefore:
We now link the geometry of the tension metric and the embedding to mechanics. The macroscopic Cauchy stress tensor of an active solid is the sum of two contributions
@@
where $p\delta_{\alpha\beta}$ is the isotropic pressure, and $\sigma_{\alpha\beta}$ is the tensile stress due to the $\uptau_{ij}$.
The Cauchy stress tensor relates the infinitesimal force $d\mathbf{f}=\bm{\sigma}\cdot\mathbf{n}\,dr$ acting on a virtual cut with the cut’s unit normal $\mathbf{n}$ and length $dr$ in the physical configuration. To see how the tension metric determines $\bm{\sigma}$, consider two adjacent cells at $\mathbf{r}$ and $\mathbf{r}+d\mathbf{r}$ (we implicitly set $dr$ to be of the scale of a single cell, infinitesimal in the continuum limit). The magnitude of the force $d\bm{f}$ acting through the virtual cut $d\mathbf{r}$ (Fig. 2) equals the tension $|d\bm{f}|=d\tau$ along the interface between the two cells. It is determined by the tension metric, $d\tau^{2}=p_{0}^{2}\,g_{\alpha\beta}(\mathbf{r})dr_{\alpha}dr_{\beta}$. On the other hand, the stress tensor applied to the cut normal yields $d\bm{f}=\bm{\sigma}\cdot\mathbf{n}\,dr=\bm{\sigma}\cdot\bm{\epsilon}\cdot d\mathbf{r}$, where $\bm{\epsilon}$ is the antisymmetric tensor and $\cdot$ is the standard Euclidean dot product. Equating tension- and cell-space expressions:
$p_{0}^{2}d\mathbf{r}^{T}\cdot\mathbf{g}(\mathbf{r})\cdot d\mathbf{r}$ $=d\tau^{2}=d\mathbf{r}^{T}\cdot(\bm{\epsilon}^{T}\!\cdot\bm{\sigma}^{2}(\mathbf{r})\cdot\bm{\epsilon})\cdot d\mathbf{r}$
$\Rightarrow\;\bm{\sigma}^{2}(\mathbf{r})$ $=p_{0}^{2}\,\bm{\epsilon}^{T}\!\cdot\mathbf{g}(\mathbf{r})\cdot\bm{\epsilon}$ (4)
Since equality holds for arbitrary $d\mathbf{r}$, the stress tensor must be the matrix square root of the metric, rotated by $\pi/2$, and scaled by $p_{0}$. We refer to Eq. (4) as the stress-metric relation. It links the microscopic stress, encoded in the tension metric, to the macroscopic stress tensor, defined in physical space. Thus, it replaces the stress-strain relation of a conventional elastic solid.
Eq. (4) depends on the cell embeddings $\mathbf{r}({\bm{\xi}})$, which are physical observables, not mere coordinates  [^2] Importantly, Eq. (4), and its consequence Eq. (20), are not “transformation laws” that relate the stress tensor in different coordinates (akin to Cauchy and Piola-Kirchhoff stress [63]). Rather, they describe how the Cauchy stress depends on deformation, like a stress-strain relation in conventional elasticity.. Of particular interest are mechanically equilibrated embeddings, which must fulfill force balance:
@@
Physically, tensile forces $\partial_{\alpha}\sigma_{\alpha\beta}$ must balance against pressure gradients $\partial_{\beta}p$. From Eqs. (4) and 5 we can already anticipate a special role for a configuration $\mathbf{r}({\bm{\xi}})$ where the tension metric is isotropic. In such an isothermal configuration, $\bm{\sigma}$ is isotropic and trivially balanced by pressure, defining an emergent stress-free reference configuration. Appendix A illustrates the link between the cell-level and the above continuum formulation for the toy problem of a uniform “rectangular foam.”
Equations (4) and (5), supplemented with mechanical boundary conditions (BCs), set the physical configuration (cell positions) $\mathbf{r}({\bm{\xi}})$ for a given tension metric $\mathbf{g}({\bm{\xi}})$ and cell density $n({\bm{\xi}})$. In the following, we introduce a mathematical framework to solve Eqs. (4)–(5) and determine when (stable) force-balanced states exist. We first study the statics problem to see how tissue shape and response to external forces emerge from fixed $\mathbf{g}({\bm{\xi}})$. Then, we turn to quasi-static shape dynamics due to changes of $\mathbf{g}({\bm{\xi}})$, and finally generalize our framework to describe plastic flow (cell rearrangement).
Figure 2: Left: In the continuum limit, the displacement vector between two cell centroids becomes a differential $d\mathbf{r}$. The tension metric determines the tension $d\tau$ on the interface between these cells. In isothermal coordinates, $dr=p_{0}^{-1}d\tau$. Right: The stress tensor $\bm{\sigma}$ is defined by the traction forces that act through infinitesimal virtual cuts. A cut along $d\mathbf{r}$ must yield the interfacial tension, relating tension metric and stress tensor in Eq. (4).
@@
To solve Eq. (5), we use complex coordinates $z=z_{1}+iz_{2}$. Complex notation converts matrix to scalar algebra and greatly simplifies the description of conformal maps, which will play an important role. A symmetric matrix $\mathbf{q}$ (e.g. the metric $\mathbf{g}$) defines a quadratic form $\mathbf{v}\mapsto q_{\alpha\beta}v^{\alpha}v^{\beta}$ which becomes
$4q_{\alpha\beta}v^{\alpha}v^{\beta}$ $=(q_{z\bar{z}}+q_{\bar{z}z})v\bar{v}+q_{zz}\bar{v}\bar{v}+q_{\bar{z}\bar{z}}vv,$ (6a)
@@
The Beltrami coefficient $\mu_{q}$ is defined by:
$\mu_{q}=\frac{q_{zz}}{q_{z\bar{z}}+\sqrt{q_{z\bar{z}}^{2}-|q_{zz}|^{2}}}=\frac{q_{zz}}{\mathrm{Tr}\,\mathbf{q}+2\sqrt{\det\mathbf{q}}}$ (7)
Since $q_{zz}$ is the deviatoric component of $\mathbf{q}$, it measures the relative anisotropy.
@@
Maps satisfying the Cauchy–Riemann equation $\bar{\partial}_{z}f=0$, i.e. $z\mapsto z^{\prime}=f(z)$, are conformal (angle-preserving). Their Jacobian is a diagonal matrix ${\mathbf{D}\kern-0.80002ptf}=\mathrm{diag}[\partial_{z}f,\overline{\partial_{z}f}]$, combining rotation and scaling by the conformal factor $\lambda_{z|z^{\prime}}=|\partial_{z}f|$. The subscript $z|z^{\prime}$ indicates the map’s source and target. Furthermore, for conformal maps, $\Delta\log\lambda_{z|z^{\prime}}=4\partial_{z}\bar{\partial}_{z}\operatorname{Re}[\log(\partial_{z}f)]=0$. More generally, the magnitude and orientation of a map’s shear is measured by the Beltrami coefficient $\mu_{z|z^{\prime}}=\bar{\partial}_{z}f/\partial_{z}f$ which vanishes for a conformal map. Note that the Beltrami coefficients of the map $f$ and of the quadratic form $({\mathbf{D}\kern-0.80002ptf})^{\dagger}\cdot{\mathbf{D}\kern-0.80002ptf}$ agree; $()^{\dagger}$ being the Hermitian adjoint. Geometrically, $f$ maps infinitesimal circles to ellipses with eccentricity $|\mu_{z|z^{\prime}}|/(1-|\mu_{z|z^{\prime}}|^{2})^{1/2}$. If $|\mu_{z|z^{\prime}}|<k<1$ everywhere, $f$ is called quasi-conformal (QC). In particular, QC maps preserve orientation. Notation-wise, their argument distinguishes QC maps $f(z,\bar{z})$ (also written $f(\mathbf{z})$), from conformal maps $f(z)$. In the following, we use the formalism of QC maps to develop our continuum theory (in Ref. [23], a similar QC approach was used to describe the growth of 2d tissues).
@@
$4g_{\alpha\beta}d\xi^{\alpha}d\xi^{\beta}$ $=(\mathrm{Tr}\sqrt{\mathbf{g}})^{2}|d\xi+\mu_{g}d\bar{\xi}|^{2}$ (9)
@@
where the Beltrami coefficient $\mu_{g}=g_{\xi\xi}/(\mathrm{Tr}[\bm{g}]+2\sqrt{\det\mathbf{g}})$ measures the anisotropy of $\mathbf{g}$. Therefore, $z(\xi,\bar{\xi})$ must fulfill a Beltrami equation:
$\bar{\partial}_{\xi}z(\xi,\bar{\xi})=\mu_{g}\partial_{\xi}z(\xi,\bar{\xi})\quad\Leftrightarrow\quad\mu_{\xi|z}=\mu_{g}$ (10)
The QC map from Lagrangian $\xi$- to isothermal $z$-coordinates exactly removes the shear part of the metric. (By abuse of notation, we will denote both the coordinates and the map $\xi\mapsto z(\xi,\bar{\xi})$ by $z$). By the Measurable Riemann Mapping theorem [25], solutions to Eq. (10) exist as long as $|\mu_{g}|<k$ for some $k<1$ (microscopically, the tension triangles must be uniformly non-degenerate). This fact will play a crucial role in guaranteeing the existence of mechanically stable states.
The metric thus becomes isotropic and is fully described by its conformal factor $\lambda_{g}$
@@
Isothermal coordinates are conformal and do not distort angles (see Fig. 3; Appendix B.2 discusses the example of a hemisphere). The Liouville equation determines the Gaussian curvature of an isotropic metric: